Write the slope-intercept equation for the line that passes through (-8, -10) and is perpendicular to -8x – 3y = -16
Answer by mananth(16949)

-8x-3y=-16
Find the slope of this line
-3y=8x+16
Divide by -3
y=-2 2/3 x-5.33
Compare this equation with y=mx+b
slope m =-2.67

The slope of a line perpendicular to the above line will be the negative reciprocal 0.38
m1*m2=-1
The slope of the required line will be 0.38 0.38 = 0.38

m=0.38 ,point(-3,-15)
Find b by plugging the values of m & the point in
y=mx+b
-15=-1.13+b
b=-13.88
m=0.38
The required equation is y=0.38x -13.88
m.ananth@hotmail.ca